# Common or specific source, features or scores; it is all a matter of information

### v1

@@ 4 Proper scoring rules
Our analysis has not made use of proper scoring rules. In [28], Vergeer ranked various “LR systems” for two hypotheses including score based methods, the common source scenario and the specific source scenario. He also argues against the claim of [20] that certain methods should not be used, albeit in a rather indirect way using (strictly) proper scoring rules ((S)PSRs) [8, 7, 6, 11, 5, 9]. Using PSRs, it is also possible to obtain the results that we derived directly in section 2.2. Although strictly speaking not necessary for this paper, we briefly discuss SPRSs for completeness.
@@
which penalises the occurrence of a hypothesis which was deemed to have small probability.
@@
The inequality (2.2) therefore can be phrased as, for $C=C_{BD}$ derived from costs $c_{ij}$,
@@
Therefore, a proof of (4.6) valid for any SPSR $C$ will imply (2.6). This is the approach taken in [28] and [5]. In fact, [28] worked in the special case of Section 2.1 (two hypotheses, two actions) and then treated several special cases that essentially differ only in the choices of $\pi$ and $e$ and then proceed in analogous ways. In [5] the general case is treated, but in a rather technical way relying on scoring rules.
@@
$-\frac{x\pi_{1}}{x\pi_{1}+\pi_{2}}\log(\frac{x\pi_{1}}{x\pi_{1}+\pi_{2}})-\frac{\pi_{2}}{x\pi_{1}+\pi_{2}}\log(\frac{\pi_{2}}{x\pi_{1}+\pi_{2}}).$
@@
For $\pi_{1}=\pi_{2}=1/2$, this quantity is called the CLLR in the literature [6, 26]. We see that this is simply the expected entropy of the posterior distribution starting with uniform priors.
5 Discussion and conclusions
The central question of this paper is whether or not additional evidence is somehow beneficial. In Sections 2.1 and 2.2 we showed that this is the case for any cost function, at least in expectation. Perhaps this sounds somewhat counter intuitive, but note that extra information can only help discovering the truth. If the cost function’s ‘goal’ is to lie about the truth, then one can tell lies better with more information.
All inequalities derived in [28] are captured by our central result (2.2). For instance, consider the situation in which one compares features with scores. As we explained in the introduction, it is not uncommon that data is summarized into scores, a process by which inevitably information is lost. Vergeer showed for this case (and, separately, for others) that on average, this lead to higher costs. (He only treated the $n=2$ case.) These conclusions all directly follow from our results. Indeed, one can interpret $\pi$ as the information one has after obtaining the scores only. They may be the joint result of an earlier prior which was updated after seeing the scores. To take the full features into account can in our set-up be interpreted as gathering extra data, and this leads therefore to lower expected costs. The same reasoning goes through for all other examples discussed in [28]. The concept of proper scoring rules, which plays an important role in his exposition is, therefore, not needed for obtaining the conclusions.
This principle also shows that calculating an LR (score-based or feature-based) is on average better than using only the prior (3.1.3 in [28]). This directly contradicts the statement made in [20] that within the Bayesian paradigm, “one cannot use score based likelihood ratios”. We analysed the arguments from [20] in some detail, as we find it important to explain how the authors of that paper arrived at their conclusion, and why this argument is incorrect.
We would like to add that the phrase ‘full-features’ is in a sense deceiving, because we do not always know what knowledge may be possible in the future. In the past, DNA identification was carried out the SGMPlus kit which contained 10 autosomal loci [18]. The more recent PPF6C kit contains 13 additional loci, so that calculating a likelihood ratio based on SGMPlus has become a score based likelihood ratio, as it comprises a data reduction relative to PPF6C. This example reveals that what we regard as a full feature method today, may be very well be a partial features method in the future, providing another argument against banning score-based methods.
In fact, every method (pathological examples aside) to calculate a likelihood ratio may in fact be regarded as score based. One method may of course be more score based than others, but even the best methods will always use some form of simplification and thus lose information that could theoretically be incorporated.
We also take a position in the alleged distinction between common source and specific source methods [21]. As we pointed out, it is a matter of available information which scenario applies, while the underlying statistical model remains the same. As with scores, we see no dichotomy between these alleged different “LR-frameworks”.
As a consequence, the question is not whether we should use score based methods, but how they can be used best. What are the conditions needed to justify data reduction, and how good are the decisions made? The answer to these questions depends entirely on the context and the objectives, see also [10]. Some score-based methods that use substantial reduction (quite naturally) perform badly [19], and may thus not be generally recommended for use in legal contexts.
The current paper, like [28] and [20], has mostly been concerned with fundamental properties of score-based methods. We hope that this work contributes in clearing score-based likelihood ratios of any fundamental suspicion, and that future research can focus on other important aspects such as calibration [32, 9, 13, 24, 25].
We close with a few remarks. First of all, one should choose the LR as corresponding to the actual knowledge. It is not allowed to compute various LRs and report the most discriminatory one.
But perhaps more importantly, when assigning a score based LR it may be the case that the evaluator had full access to all obtained data, but fed it into an algorithm for computing, say a common source LR. Then, it may well be the case that, even if no statistical model is available for the full data, common sense of the evaluator may lead them to assess globally in which direction the LR would likely move if that model were available. We give an example, computed with the previously discussed toy model of [20]. Suppose that $e=(\theta_{x},(e_{u},e_{s}))=(0,(5,5))$. This yields $LR_{SS}(e)=LR(\theta_{x},(e_{u},e_{s}))=0.0134$ and $LR_{CS}(e)=LR(e_{u},e_{s})=5.555.$ In this case, if we only process the two measurements, since both traces have the same value, the common source LR is large. If we evaluate the specific source LR, then the observation $e_{s}$ is redundant and we need to obtain measurement 5 from a source with parameter 0, or from a random source. Finding $e_{x}=5$ is very unlikely if $\theta_{y}=0$ and $\sigma_{x}^{2}=2$, compared to obtaining it from a random source.
If we would not have the model for $LR_{SS}$, but only for $LR_{CS}$, we could nevertheless have an idea of the distribution of sources, even if not specified in a probability distribution. In that case, care should be taken, since it may be clear to the researcher that the computed $LR_{CS}$ is likely larger than the LR that would be obtained from the full data. But this is matter of caution that applies to almost any statistical model: at all times, one must have a justification of why this model would apply in the situation at hand, and why, if more data are available to the evaluator than those summarized into a LR, the computed LR can be taken as an accurate or conservative assessment of the whole data.
A similar situation has long existed for the interpretation of DNA mixture profiles: models that gave a LR based on the recorded alleles were available, but these models could not process peak height data. Thus, instead of the full data, only a part was processed in the resulting LR, but the forensic analysts and interpretators had access to the whole profile. They would then, visually and qualitatively and not leading to additional LR assessment, inspect the correspondence between PoI and trace profile to assess whether it is reasonable to expect, according to their expertise, that the LR would not decrease when peak heights would, additionally, be taken into account. Of course, if it would be discovered that the peak heights were inconsistent with contribution, the computed LR would be overruled. This way, the full data were used qualitatively only.
It is of course easy to conceive, given the profile of a PoI and observed alleles of a trace profile so that a large LR for contribution results, of peak heights that are inconsistent with the PoI’s contribution and thus would reduce the LR to zero. However, this is mitigated in practice by the fact that, when strong evidence is obtained, this usually means that the posterior probability of contribution is large, and for actual contributors peak heights speaking against contribution are rare and the changes in LRs they would cause are bounded by (2.9).
Thus, we do not believe that any LR model should be discredited on the basis of the information that it does or does not process. It may in fact be preferable to have a more accurate model for a simplification of the data, than a flawed model for the full data. This can perfectly go together with the evaluator also not being discharged of trying to interpret all relevant data they have access to, even if only qualitatively.

### v2

@@ 4 Discussion and conclusions
Our analysis has not made use of proper scoring rules, but for completeness we discuss some of the relevant aspects in this appendix. In vergeer, Vergeer ranked various “LR systems” for two hypotheses including score based methods, the common source scenario and the specific source scenario. He also argues against the claim of neumann that certain methods should not be used, albeit in a rather indirect way using (strictly) proper scoring rules ((S)PSRs) degroot; dawid2007; brummer2006; raftery; brummer; ferrer. Using PSRs, it is also possible to obtain the results that we derived directly in section 2.2. Although strictly speaking not necessary for this paper, we briefly discuss SPRSs for completeness.
@@
which penalizes the occurrence of a hypothesis which was deemed to have small probability.
@@
The inequality in Theorem 2.1 therefore can be phrased as, for $C=C_{BD}$ derived from costs $c_{ij}$,
@@
Therefore, a proof of (4.6) valid for any SPSR $C$ will imply Theorem 2.1. This is the approach taken in vergeer and brummer. In fact, vergeer worked in the special case of Section 2.1 (two hypotheses, two actions) and then treated several special cases that essentially differ only in the choices of $\pi$ and $e$ and then proceed in analogous ways. In brummer the general case is treated, but in a rather technical way relying on scoring rules.
@@
$\frac{x\pi_{1}}{x\pi_{1}+\pi_{2}}\log(\frac{x\pi_{1}}{x\pi_{1}+\pi_{2}})-\frac{\pi_{2}}{x\pi_{1}+\pi_{2}}\log(\frac{\pi_{2}}{x\pi_{1}+\pi_{2}}).$
@@
For $\pi_{1}=\pi_{2}=1/2$, this quantity is called the CLLR in the literature brummer2006; vanlierop2024. We see that this is simply the expected entropy of the posterior distribution starting with uniform priors.